# The Noisy Path from Source to Citation: Measuring How Scholars Engage with Past Research

### v2

@@ 5 Higher Fidelity is Associated with Closer Proximity
First, our findings reveal that proximity—whether in terms of access, disciplinary context, temporal closeness, or author familiarity—plays a critical role in shaping citation fidelity. Self-citations demonstrate higher fidelity scores than non-self-citations. Researchers citing their own work are naturally more familiar with the material and better positioned to reproduce the original context and content. This phenomenon serves as a double validation for our citation fidelity measure, as it reflects the expected outcome: authors engaging more deeply with their own work and citing with fidelity.
Similarly, open-access publications exhibit higher fidelity compared to non-open-access publications, likely because unrestricted access to the original sources allows more direct engagement and accurate representation of cited work. The absence of access barriers facilitates authors’ ability to thoroughly examine and interpret the cited material, thus reducing potential information loss Nagaraj et al. (2020). Citation fidelity is higher for within-field citations compared to cross-field citations. This is likely because authors are more familiar with the methodologies, terminologies, or conceptual frameworks of their own disciplines, and are able to represent without much information change. In contrast, cross-field citations, while essential for interdisciplinary knowledge transfer Thilakaratne et al. (2018), may introduce additional information loss due to differences in their research norms, jargon, or methodologies.
Second, our findings suggest that the roles and collaborative dynamics within teams can shape how prior work is represented. Interestingly, citation fidelity decreases as the seniority of first authors increases, while showing no significant relationship with the seniority of last authors—the traditional role for team leaders or principal investigators. This distinction may reflect the different roles and responsibilities within team collaboration. Last authorship often involves advisory or supervisory duties, where responsibility for citation accuracy is typically delegated to other team members Brand et al. (2015); Sauermann and Haeussler (2017).
The decline in citation fidelity with increasing first-author seniority may also indicate a shift in focus among senior researchers. Senior authors may rely more on their established reputations or heuristics, potentially leading to less detailed engagement with individual citations. This highlights how differences in team roles and levels of engagement with cited works can influence the accuracy of scholarly representation.
@@
A key concern in citation practices is that authors may not always engage directly with the sources they cite Kaplan (1965). Instead, they often rely on intermediary sources—papers that have already interpreted and cited the original work. This intermediary engagement may introduce misinterpretations or distortions, which may propagate across the literature through citation chains Simkin and Roychowdhury (2002); West and Bergstrom (2021). Much like the “telephone game,” where a message is whispered from person to person and gradually distorted, scientific claims can evolve—or degrade—as they are repeatedly paraphrased, summarized, or reinterpreted in subsequent research. We employ a quasi-causal experiment setting to examine how intermediary engagement with prior work impacts knowledge transmission, as reflected in their citation fidelity.
@@
Experimental setup To test these hypotheses, we conduct a quasi-experimental study, identifying approximately 50K paper pairs (C and D) with the following setup: both papers cite the same claim from same original paper A, but paper D cites A directly (control group), while paper C cites both A and an intermediary paper B, which itself cites A (treatment group). To control for potential confounding factors, papers C and D are exactly matched by publication year and field, for comparability across both groups.
We use citations to an intermediary paper as a proxy for engagement with that paper. While this does not guarantee direct influence or even that the intermediary paper was read, it reflects a higher likelihood of engagement. Consequently, citing an intermediary paper increases the likelihood of being influenced by its interpretation of the original claims compared to papers that do not cite any intermediary source.
Results In the treatment group (citing both paper A and paper B), citation fidelity to paper A is 0.06 lower than in the control group (citing only paper A). This finding supports H1, indicating that engagement with intermediary sources may contribute to low citation fidelity. The fidelity of paper B’s citation to paper A correlates with the fidelity of paper C’s citation to paper A, corroborating H2. Low citation fidelity in the intermediary source can lead to further information loss in subsequent citations.
These results show the inherent risks associated with relying on intermediary sources in scholarly communication. While intermediaries often serve as practical and time-efficient tools for navigating the vast body of scientific literature, they can also become vectors for information loss or even misrepresentation Sharkey et al. (2023). This is particularly concerning when intermediaries fail to accurately reflect the original claims, as their interpretations may cascade through subsequent citations, compounding distortions over time West and Bergstrom (2021). Such reliance not only affects the fidelity of knowledge transmission but also raises broader questions about the robustness of scholarly practices. Future work could investigate the specific mechanisms through which these distortions occur, such as affirmative citation or selective reporting, to develop targeted interventions for improving citation fidelity.
@@ 7 Conclusion
In this study, we introduced a computational pipeline to analyze citation fidelity on a large scale and developed an automated measure to evaluate how authors report prior findings. By applying this method to 13 million citation sentence pairs, we uncovered several key insights into the dynamics of scholarly communication and show citation fidelity is higher with closer proximity. Through a quasi-causal experiment, we further establish that the “telephone effect” contributes to information loss: when citing papers exhibit low fidelity to the original claim, future papers that cite both the citing paper and the original show even lower fidelity to the original.
This study addresses two critical research gaps: the lack of large-scale measures to assess citation fidelity and the absence of empirical evidence on how authors engage with prior literature. By operationalizing citation fidelity and exploring its contextual drivers, we reveal systematic differences in citation fidelity and provide a more nuanced understanding of how information is preserved or lost during the citation process. These findings not only provide evidence on the mechanisms of knowledge transmission within the scientific community but also reveal the limitations of traditional citation analyses, which often overlook dimensions of fidelity and engagement.
Finally, this study represents an initial step toward measuring a critical source of distortion in scientific evidence. This work contributes to broader discussions on the reliability of citations as indicators of research quality and influence, as well as ongoing efforts to address misinformation in science and questionable responsible research practices.
@@ 8 Limitations
While citation fidelity scores capture alignment across several dimensions—such as topical alignment, generality, and accurate representation of the original claim—they reflect an aggregate measure rather than disentangling specific components like phrasing accuracy or conceptual alignment. Consequently, these scores may oversimplify nuanced aspects of citation behavior, as they do not directly correspond to the widely discussed concept of a “miscitation,” often characterized by the propagation of incorrect or misleading information.
This measure may be influenced by factors beyond the scope of this study, such as varying levels of generalization, synthesis with other sources, or factual inaccuracies in reported claims. While we consider this measure a meaningful first step toward understanding citation fidelity, it does not fully address the broader challenges of misinformation in citations or questionable research practices. Future research could explore these nuanced aspects to gain a deeper understanding of how and why information fidelity deteriorates during the citation process.
@@ Appendix A Appendix
By filtering for sentences dedicated to reporting findings from a single source, we create a more controlled and precise dataset for analyzing citation fidelity. This structured approach helps isolate the core reporting function of citations, minimizing confounding factors and improving the reliability of our analysis. By filtering for sentences dedicated to reporting findings from a single source, we create a controlled dataset that isolates the core reporting function of citations,improving the reliability of our analysis.
@@ Appendix B Regression Analysis Details and Results
Figure 7 presents the regression coefficients for additional factors (beyond those discussed in Section 5) affecting citation fidelity, with 95% confidence intervals. Categories within predictors (e.g., field, publication type) are color-coded for clarity. The reference categories for categorical predictors are as follows: “Field: Physics,” “OpenAccess: False,” “Publication Type: Other,” “Publication Year: 2000,” “Within-Field: False,” and “Self-Cited: False.” For continuous variables, they are grouped into bins for visualization purposes to better illustrate trends (in these cases, the bin corresponding to the smallest value serves as the reference category).
@@
Citation fidelity decreases as the citation count of the cited paper increases. This trend suggests that heavily cited papers are often engaged with less deeply, possibly due to their canonical status or the reliance on large number of secondary interpretations. Papers with high citation counts may be cited symbolically or to signal familiarity rather than as direct sources of content, leading to a low fidelity their representation.
Citation fidelity also increase with longer citation contexts. However, the increase diminish beyond approximately 100 words, indicating that longer contexts are associated with more thorough explanations of the cited work, but excessively long contexts may dilute focus or introduce irrelevant details. Similarly, frequent citation of a source within a paper is associated with lower fidelity, suggesting more extensive interpretation and paraphrasing of the cited content. When a source is referenced multiple times, authors may engage with different aspects of the cited work rather than reiterating the same claim, leading to greater variation in fidelity.
Figure 7: Regression coefficients with 95% confidence intervals for predictors of citation fidelity. Each point represents the estimated effect of a predictor (e.g., discipline, publication type, context length) on citation fidelity. Horizontal lines indicate the 95% confidence intervals for each coefficient. Statistically significance level is indicated by asterisks (*** for p < 0.001).

### v3

@@ Results
Second, our findings suggest that the roles and collaborative dynamics within teams can shape how prior work is represented. Interestingly, citation fidelity decreases as the seniority of first authors increases, while showing no significant relationship with the seniority of last authors—the traditional role for team leaders or principal investigators. This distinction may reflect the different roles and responsibilities within team collaboration. Last authorship often involves advisory or supervisory duties, where responsibility for writing detailed citations is typically delegated to other team members, usually the first author. Brand et al. (2015); Sauermann and Haeussler (2017).
The decline in citation fidelity with increasing first-author seniority may indicate a shift in how senior researchers approach citations. More experienced authors might rely more on their established heuristics and prior knowledge, or adopt a more top-down writing approach, resulting in less engagement with the cited materials.
@@
A key concern in citation practices is that authors may not always engage directly with the sources they cite Kaplan (1965). Instead, they often rely on intermediary sources—papers that have already interpreted and cited the original work. This intermediary engagement may introduce misinterpretations or distortions, which may propagate across the literature through citation chains Simkin and Roychowdhury (2002); West and Bergstrom (2021). Much like the “telephone game,” where a message is whispered from person to person and gradually distorted, scientific claims can evolve or degrade, as they are repeatedly paraphrased, summarized, or reinterpreted in subsequent research. We employ a quasi-causal experiment setting to examine how intermediary engagement with prior work impacts knowledge transmission, as reflected in their citation fidelity.
@@
Experimental setup
To test these hypotheses, we conduct a quasi-experimental study, identifying approximately 50K paper pairs (C and D) with the following setup: both papers cite the same claim from same original paper A, but paper D cites A directly (control group), while paper C cites both A and an intermediary paper B, which itself cites A (treatment group). To control for potential confounding factors, papers C and D are exactly matched by publication year and field, for comparability across both groups.
We use citations to an intermediary paper as a proxy for engagement with that paper. While this does not guarantee direct influence or even that the intermediary paper was read, it is associated with a higher chance of engagement. Consequently, citing an intermediary paper increases the likelihood of being influenced by its interpretation of the original claims compared to papers that do not cite any intermediary source.
Results
In the treatment group (citing both paper A and paper B), citation fidelity to paper A is 0.06 lower than in the control group (citing only paper A). This finding supports H1, indicating that engagement with intermediary sources may contribute to low citation fidelity.
We further divide the treatment group based on the fidelity of the intermediary citation. Specifically, the fidelity of paper B’s citation to paper A is classified into three levels: high (score above 4), medium (scores of 3–4), and low (score below 3). We find the fidelity of paper C’s citation to paper A positively correlates with the fidelity of paper B’s citation to paper A, corroborating H2. Low citation fidelity in the intermediary source can lead to further information loss in subsequent citations.
These results show the inherent risks associated with relying on intermediary sources in scholarly communication. While intermediaries often serve as practical and time-efficient tools for navigating the large body of scientific literature, they can also become vectors for information loss or even misrepresentation Sharkey et al. (2023). This is particularly concerning when intermediaries fail to accurately reflect the original claims, as their interpretations may cascade through subsequent citations, compounding distortions over time West and Bergstrom (2021). Such reliance not only affects the fidelity of knowledge transmission but also raises broader questions about the robustness of scholarly practices.
Our quasi-experimental results reflect the average effect of having any intermediary source on citation fidelity. However, intermediary effects are likely more nuanced. The number of intermediaries involved, as well as their characteristics, such as recency, prestige, and citation count, may differentially influence how interpretations are transmitted. Future work could explore these dimensions to better understand the role intermediary sources play and investigate the specific mechanisms through which these low fidelity occur, such as affirmative citation or selective reporting to understand the source of low fidelity.
@@ 7 Conclusion
In this study, we introduced a computational pipeline to analyze citation fidelity on a large scale and developed an automated measure to evaluate how authors report prior findings. By applying this method to 13 million citation sentence pairs, we uncovered several key insights into the dynamics of scholarly communication and showed citation fidelity is higher with closer proximity. Through a quasi-causal experiment, we further establish that the “telephone effect” contributes to information loss: when citing papers exhibit low fidelity to the original claim, future papers that cite both the intermediary and the original show even lower fidelity to the original.
This study addresses two critical research gaps: the lack of large-scale measurements to assess citation fidelity and the lack of empirical evidence on how authors engage with prior literature. By operationalizing citation fidelity and exploring its contextual drivers, we reveal systematic differences in citation fidelity and provide a more nuanced understanding of how information is preserved or lost during the citation process. These findings not only provide evidence on the mechanisms of knowledge transmission but also reveal the limitations of traditional citation analyses which often overlook dimensions of fidelity and authors’ engagement.
Finally, this study represents an initial step towards enriching citation analysis with text-level nuance. It could help advance our understanding of citation quality—as a potential source of distorted or misrepresented information, as a behavioral trace of author engagement with prior work, and as a valuable complement to existing metrics for research evaluation.
@@ 8 Limitations
Citations are made for different purposes and with different functions, and authors employ a variety of ways in writing to paraphrase, summarize, or reinterpret prior work. This variety makes systematically measuring citation fidelity challenging. In our analysis, we apply a series of filters to both citing and cited sentences to form appropriate citation pair instances—for example, limiting citing sentences to those with a single source cited at the end, and limiting cited sentences to single-sentence claims labeled as results or conclusions. These filters help us isolate citation pairs where claims are expected to be reported faithfully. They prioritize precision but at the expense of recall and they may also introduce potential selection bias and overlook alternative citation forms that are also interesting to investigate.
In particular, we restrict the analysis to one-to-one, sentence-level matches between citing sentences and candidate claims in the cited paper, with the fidelity score proxied as the highest alignment of each sentence pair. However, in practice, authors may distribute their reporting text across multiple sentences or use only a brief clause. Likewise, cited claims may also be diffuse, spanning several segments of text in the cited work. As a result, this simplification could likely underestimate fidelity in more complex rhetorical structures. Future work could improve upon this by expanding the scope of citation context and claim spans.
Also, while our citation fidelity scores capture alignment across several dimensions—such as topical alignment, generality, and factual representation of the original claim—they reflect an aggregate measure rather than disentangling these specific components. So these scores may oversimplify nuanced aspects of citation behavior and may be influenced by factors such as varying levels of abstraction, synthesis with other sources, or factual inaccuracies in reported claims. These factors may blur the boundary between acceptable paraphrasing and substantive misrepresentation. Therefore, while our fidelity measure represents a meaningful first step toward identifying potential sources of misrepresented scientific evidence, it does not correspond directly to the widely discussed concept of “miscitation” or “reference error.” Future research should explore these finer-grained aspects, not only to identify such cases but also to gain deeper insights into how and why misinformation or questionable research practices emerge and propagate through the citation process.
9 Acknowledgments
We thank members of Blablablab and DiscoveryLab, as well as participants in the Computational Social Science Working Group at the University of Michigan, for their insightful feedback and suggestions on this project.
@@ Appendix A Appendix
By filtering for sentences dedicated to reporting findings from a single source, we create a controlled and precise dataset that isolates the core reporting function of citations, improving the reliability of our analysis.
@@ Appendix B Regression Analysis Details and Results
Figure 7 presents the regression coefficients for additional factors (beyond those discussed in Section 5) affecting citation fidelity, with 95% confidence intervals. Categories within predictors (e.g., Field.of.Study and Publication.Type) are color-coded for clarity. The reference categories for categorical predictors are as follows: “Field: Physics,” “OpenAccess: False,” “Publication Type: Other,” “Publication Year: 2000,” “Within-Field: False,” and “Self-Cited: False.” For continuous variables, they are grouped into bins for visualization purposes to better illustrate trends (in these cases, the bin corresponding to the smallest value serves as the reference category).
@@
Citation fidelity decreases as the citation count of the cited paper increases. This trend suggests that heavily cited papers are often engaged with less deeply, possibly due to their canonical status or the reliance on a large number of secondary interpretations. Papers with high citation counts may be cited symbolically or to signal familiarity rather than as direct sources of content, leading to a low fidelity in their representation.
Citation fidelity also increases with longer citation contexts. However, the increase diminishes beyond approximately 100 words, indicating that longer contexts are associated with more thorough explanations of the cited work, but excessively long contexts may dilute focus or introduce irrelevant details. Similarly, frequent citation of a source within a paper is associated with lower fidelity, suggesting more extensive interpretation and paraphrasing of the cited content. When a source is referenced multiple times, authors may engage with different aspects of the cited work rather than reiterating the same claim, leading to greater variation in fidelity.
Figure 7: Regression coefficients with 95% confidence intervals for predictors of citation fidelity. Each point represents the estimated effect of a predictor (e.g., discipline, publication type, context length) on citation fidelity. Horizontal lines indicate the 95% confidence intervals for each coefficient. The statistical significance level is indicated by asterisks (*** for p < 0.001).